OpenStudy (anonymous):

Solve 2x+9(x^1/2)-5 Please show steps!

OpenStudy (anonymous):

\[= 2x + 9x ^{1.5} - 5\] ok?

OpenStudy (anonymous):

@Kenneditodd what does the question want? the mesure of x?

OpenStudy (anonymous):

That means the square root of x....the equation is equal to zero, and yes it wants x.....I know the answer I just need to know how to get it

OpenStudy (anonymous):

\[x^{1.5} = \sqrt{x ^{3}}\] ok?

OpenStudy (anonymous):

@Kenneditodd , OK?

OpenStudy (anonymous):

No it's x=1/4

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

ok ok

OpenStudy (anonymous):

but I didn't find the answer...

OpenStudy (anonymous):

Yes....and I think you misunderstood the problem, it's 2x+9 times the square root of x -5 @PFEH.1999

OpenStudy (anonymous):

X-5 is not a binomial btw

OpenStudy (anonymous):

@sweetsandeverything yes

OpenStudy (anonymous):

If we are looking for the zero of the expression, we can make a substitution to make it easier. Let u = sqrt x, then\[0=u^2+9u-5=(2u-1)(u+5) \implies u=\frac{1}{2}= \sqrt {x} ~or~u=-5\]\[\implies x=\frac{1}{4}\]The solution u=-5 is discarded, since the square root is positive.

OpenStudy (mathstudent55):

\(2x + 9\sqrt x - 5 = 0\) Let \(u = \sqrt x \), then \(u^2 = x\) Substitute in original eqution: \(2u^2 + 9u - 5 =0 \) \((2u - 1)(u + 5) =0 \) \(2u - 1 = 0\) or \(u + 5 = 0\) \(2u = 1\) or \(u = -5\) \(u = \dfrac{1}{2} \) or \(u = -5\) Since x = u^2, \(x = \left( \dfrac{1}{2} \right)^2 \) \(x = \dfrac{1}{4} \) or \(x = (-5)^2 \) \(x = 25\) When you check the two solutions in the original eqaution, x = 25 does not work, so the solutioon is \(x = \dfrac{1}{4} \)